OpenStudy (anonymous):

Solve 4h = 92 for h. H=

OpenStudy (anonymous):

divide both sides by 4

OpenStudy (awy):

H=23

OpenStudy (esshotwired):

to solve for h you must get it alone. to do so you must divide both sides by 4 to cancel out the 4 with the h. then you will get h=23. does that make more sense to you?

OpenStudy (anonymous):

Isolate h, to do so divide (the inverse of multiplication. 4h = 92 /4 /4 h = 23

OpenStudy (anonymous):

That's it! :D

OpenStudy (anonymous):

h=23

OpenStudy (anonymous):

You can check, 4*23 does = 92 :)

OpenStudy (anonymous):

what is it 92 or 23

OpenStudy (anonymous):

23

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4*23 = 92 ._.

OpenStudy (anonymous):

Divide both sides by 4